# UTMath: Math Evaluation with Unit Test via Reasoning-to-Coding Thoughts

### v1

@@ Abstract
The evaluation of mathematical reasoning capabilities is essential for advancing Artificial General Intelligence (AGI). While Large Language Models (LLMs) have shown impressive performance in solving mathematical problems, existing benchmarks such as GSM8K and MATH present limitations, including narrow problem definitions with specific numbers and reliance on predetermined rules that hinder accurate assessments of reasoning and adaptability. This paper introduces the UTMath Benchmark, which robustly evaluates the models through extensive unit tests. It consists of 1,053 problems across 9 mathematical domains, with over 68 test cases per problem. We propose an innovative evaluation framework inspired by unit testing in software development, focusing on both accuracy and reliability of results. Furthermore, we introduce the Reasoning-to-Coding of Thoughts (RCoT) approach, which encourages LLMs to perform explicit reasoning before generating code, leading to generating more advanced solution and improved performance. Furthermore, we are releasing not only the UTMath benchmark but also the UTMath-Train training dataset (more than 70k samples), to support the community in further exploring mathematical reasoning.
@@ 1 Introduction
Figure 1: Comparing UTMath with MATH. In the left, GPT-4o easily solved a question but failed on a similar question with a different number. The right figure illustrates our benchmark, where each problem includes multiple test cases, and a solution must pass all test cases to be strictly considered correct. We also propose a new prompting method RCoT, where the LLM first reasons through the problem and then generates code.
@@
However, classic benchmarks exhibit several limitations that impede the accurate and comprehensive assessment of these models’ capabilities Ahn et al. (2024). First, these benchmarks test models on narrowly defined problems with some specific numbers, which do not adequately assess adaptability to similar but varied scenarios as shown in Fig. 1. Second, their evaluation relies on predetermined rules or the method of LLM-as-a-Judge( Dubois et al. (2024); Zheng et al. (2023)) that usually failed with capricious responses of LLMs. For example, an accurate answer need to be extracted to exactly match the fianl answer in the dataset GSM8K, TheoremQA, and MATH dataset. Furthermore, these benchmarks focus more on final answers than on the underlying reasoning steps. While recent work has made great progress in developing new benchmarks for evaluating the mathematical reasoning of LLMs, many of these approaches still fall short of addressing the fundamental limitations of earlier datasets. For instance, benchmarks like GSM-HARD Gao et al. (2023), GSM-IC Shi et al. (2023), GSM-Plus Li et al. (2024a), MetaMath Yu et al. (2023) have extended the dataset of GSM8K or MATH with some perturbation such as substitution, reversing, distractor insertion. These efforts, while valuable, are characterized by limited coverage and high costs.
@@
In this context, our work seeks to bridge these gaps by proposing a solid and robust benchmark that accurately evaluates the mathematical capabilities of LLMs.Drawing inspiration from evaluation methods in software development, where solution correctness is determined through comprehensive unit testing, a solution is deemed correct if it passes all tests, as this demonstrates that its logic is sufficiently robust to generalize across a broad spectrum of scenarios. Similarly, we propose designing a comprehensive set of unit tests for mathematical problems to rigorously assess the reasoning processes of LLMs. If a solution passes all unit tests across a group of similar problems, it suggests that the reasoning behind the solution is more reliable and trustworthy. Specifically, we introduce the UTMath, a novel benchmark derived from the On-Line Encyclopedia of Integer Sequences (OEIS) OEIS Foundation Inc. (2024). The benchmark consists of 1,053 problems spanning 9 mathematical domains, such as Number Theory and Geometry. Each problem is accompanied by more than 68 test cases that provide a set of inputs and their corresponding outputs.
@@
In terms of evaluation methodology, our benchmark requires models to derive a general solution for a class of problems, typically represented in the form of code. Compared to solving a problem defined by specific numbers, developing such a general solution is substantially more challenging, requiring higher levels of intelligence and reasoning ability. However, we observed that when the model is required to perform reasoning and coding in a single response, as in “Program of Thoughts (PoT)” Chen et al. (2022), it consistently tends to produce simpler, more straightforward solutions. We surmise that this tendency may be influenced by the distribution of coding data. To address this, we introduced the “Reasoning-to-Coding of Thoughts (RCoT)” approach, which requires the LLM to perform mathematical reasoning in the first turn without any coding instruction, then writing code based on that reasoning. Compared to PoT, RCoT shifts the code distribution towards mathematics in the first turn, prompting more reasoning steps.
@@
We conducted a comprehensive study with 8 LLMs. Some of our key findings are summarized as follows: (1) The best model, GPT-4o, only solves 26.93% problem in our benchmark, demonstrate the difficulty of our benchmarks. (2) Modern LLMs perform poorly in Graph Theory, Group Theory, Geometry and Topology (Fig. 4). (2) With RCoT, 7 of 8 evaluated LLMs generated more efficient solutions, with most models achieving higher scores (Fig. 2). (3) RCoT can significantly improve the pass@k performance of LLMs (§ 5.4). (4) The quality of reasoning significantly impacts the accuracy and efficiency of the model’s final solution (§ 5.5). More interesting findings can be found in § 5. We hope our findings contribute to a deeper understanding of current reasoning ability of LLMs and the further development of models.
@@ 2.1 Benchmarks
Besides, researchers have also introduced many others focused on evaluating various aspects of LLMs like the robustness. These include GSM8K-based variants: GSM-8K-Adv  Anantheswaran et al. (2024), GSM-Hard  Gao et al. (2023), GSM-Plus  Li et al. (2024a), GSM-IC  Shi et al. (2023), and several independent benchmarks: Mathattack  Zhou et al. (2024b), MetaMath  Yu et al. (2023), PROBLEMATHIC  Anantheswaran et al. (2024), MATHCHECK  Zhou et al. (2024a), as well as other benchmarks  Li et al. (2024b, 2023).
The distinctions between our proposed benchmark and existing ones are as follows:
$\bullet$ Multiple Case Validation Instead of using single cases that can be memorized, our questions are sequence-based, allowing numerous cases for validating true understanding.
$\bullet$ True Reasoning Evaluation Hard cases and runtime metrics help filter memorization and compare solution efficiency, precisely assessing reasoning abilities.
@@ 2.3 Prompting Methods
The Chain-of-Thought Wei et al. (2022) prompting technique encourages models to express reasoning steps in natural language before concluding. Similarly, the approach by  Kojima et al. (2022) uses the phrase "Let’s think step by step" to effectively guide large language models through their reasoning.
Inspired by CoT, several effective prompting methods have been developed, such as Tree-of-Thoughts Yao et al. (2024), Graph-of-Thoughts Besta et al. (2024).
Program-of-Thought prompting Chen et al. (2022): PoT generates programs as the intermediate steps and integrates external tools like a Python interpreter for precise calculations, as well as other prompting methods Wang et al. (2023); Gao et al. (2023).
Our Reasoning-Coding of Thoughts (RCoT) method stands out by dividing reasoning into two steps: reasoning and implementing based on reasoning. This segmentation provides deeper insights into LLM reasoning abilities. Coding reduces errors from limited computational skills and evaluates solutions for better success and efficiency. The advantages can be summarized as follows:
$\bullet$ Code Output Evaluation We require LLMs to output code, focusing on reasoning rather than direct answers, to better reflect their reasoning skills.
$\bullet$ Observation of Reasoning Process By mandating code implementation, we can observe and validate the LLM’s reasoning process, not just the final answer.
@@ 3 UTMath Benchmark
The OEIS was established to document integer sequences of interest to both professional and amateur mathematicians, and it has become widely cited in the mathematical community. As of February 2024, it contains over 370,000 sequences OEIS Foundation Inc. (2024). Each sequence is accompanied by an identification number, a brief description, some sample integers, generation rules, links to relevant literature, and, where possible, program code for computing the sequences. An example sequence is shown in Apendix A.
3.2 Benchmark Construction
The sequences that meet our expectations and can be used to test reasoning abilities should satisfy the following conditions:
  1. 1.
    The sequence should require reasoning to derive its recurrence or formula, not just simple calculations.
  2. 2.
    The sequence must be derivable with infinitely many terms, allowing deduction of a(n) for any positive integer n.
  3. 3.
    The sequence must be accurately described to deduce it entirely from the description.
  4. 4.
    The sequence should belong to mathematics, preventing reasoning errors from lack of domain knowledge.
Data Crawling
@@ 3.2 Benchmark Construction
OEIS provides users with a list of principal sequences  [^1] https://oeis.org/wiki/Index_to_OEIS, which are representative sequences in the OEIS. To help users quickly find the main sequences within their areas of interest, OEIS categorizes these sequences into 118 sections based on the first 2-3 letters of their content themes. By scraping the category tags within each section and the AIDs of their subordinate sequences, we obtained 569 categories and 23,238 principal sequences’ AIDs. OEIS provides an interface to request the JSON data of the HTML page for each sequence using its AID  [^2] https://oeis.org/wiki/JSON_Format. By passing the sequence AIDs to this interface, we acquired the JSON data for these 23,238 sequences.
Data Cleaning
@@
We found that some of the sequences we collected did not meet our criteria and should be removed. Here, we present only a portion, with further details provided in the appendix  B.
@@
$\bullet$ Hard to solve, few terms are discoverable A portion of the sequences retrieved are marked as “hard” in the keyword field of their entries in OEIS. According to OEIS, “Any sequence which can be extended only by new ideas, rather than more computation deserves keyword: hard. Similarly, if computing a term of the sequence would probably merit a paper in a peer-reviewed journal (discussing the result, the algorithm, etc.)”  [^3] https://oeis.org/wiki/User:Charles_R_Greathouse_IV/Keywords/difficulty Another related keyword attribute is "fin" (finite), indicating sequences with limited length. For our purposes, an "appropriate" sequence should possess an infinite number of terms and be infinitely derivable.
@@
$\bullet$ Difficult to Generate Programmatically In OEIS, most sequences are provided with fields such as Mathematica, program, or formula, but not all sequences include these fields. Since OEIS does not specify the reasons for missing these fields, we assume that these sequences might be difficult to generate programmatically.
@@
$\bullet$ Overly Simple Some sequences are simple, needing little inference with clear solving algorithms. Due to uncertainty in sequence description, we use GPT-4o to determine if a sequence requires reasoning or just implementation; if mostly implementation, it’s excluded. For example, A000178444https://oeis.org/A000178: "Superfactorials: product of first n factorials.", which is a sequence that merely requires implementation will be removed.
After addressing the aforementioned issues, we ultimately obtained 1053 sequences that meet our criteria.
Standardization of Question Statements
As OEIS is a specialized academic tool, we cannot directly use the descriptions provided for each sequence as problem statements. This is primarily due to the following reasons:
$\bullet$ Specialized Terminology Some sequence descriptions use complex math terms that need examples or explanations to be clear. Using them directly as problems might test mathematical knowledge rather than reasoning skills. So, it’s important to explain key concepts to focus on reasoning and reduce the extra knowledge needed.
$\bullet$ Brevity and Ambiguity Some sequence descriptions are excessively brief and lack a clear definition of what a(n) is.
We used GPT-4o to standardize these by adding background info without revealing solutions and making the language smoother. More details are in the appendix 10. Despite our requirement that ensuring the consistency of meaning between the original and processed descriptions, hallucinations can still occur. To mitigate this, we performed a manual verification of the standardized problem statements to ensure they matched the original descriptions in meaning and were easy to comprehend.
Figure 3: Comparison between original and standardized problem statements. The standardized version includes hints and explains the specific meaning of $\sigma(x)$.
Hard Test Cases Mining
@@
The primary goal of this paper is to evaluate the reasoning ability of LLMs. Generally, more efficient solutions to a problem imply stronger reasoning capabilities. Therefore, we aim for our evaluation to distinguish whether a solution is efficient. However, in the OEIS (Online Encyclopedia of Integer Sequences), each sequence only lists the first few n terms, normally n<100, which can be obtained without requiring particularly efficient methods. This limitation prevents the evaluation from effectively distinguishing between efficient and inefficient solutions. An obvious fact is that the difficulty of computing the first 10 terms of a sequence within a time limit is significantly different from computing terms starting from the $10^{6}$th term. Therefore, we aim to create more challenging test data to better assess the reasoning capabilities of LLMs.
@@
Fortunately, many OEIS sequences include corresponding Mathematica code that can be regarded as the ground-truth solution for each problem. We extract this Mathematica code for each sequence, formalizing it to compute the first $N$ terms, $A_{1},...,A_{N}$, of the sequence. Using a binary search algorithm, we determine the maximum value of $N_{max}$ for which the code can compute the sequence within 10 seconds, where we set $10^{6}$ as the upper bound. Finally, we add the last 10 terms $A_{N_{max-9}},...,A_{N_{max}}$ into our benchmark as the hard test cases to evaluate the complexity of a solution. Our experiments demonstrate that these cases precisely differentiate more efficient and intelligent solutions.

### v2

@@ Abstract
The evaluation of mathematical reasoning capabilities is essential for advancing Artificial General Intelligence (AGI). While Large Language Models (LLMs) have shown impressive performance in solving mathematical problems, existing benchmarks such as GSM8K and MATH present limitations, including narrow problem definitions with specific numbers and reliance on predetermined rules that hinder accurate assessments of reasoning and generality. This paper introduces the UTMath Benchmark, a robust evaluation framework designed to assess LLMs through extensive unit tests, with a focus on both the accuracy and generality of model responses. It comprises 1,053 cutting-edge problems spanning nine mathematical domains, with an average of 68 test cases per problem. UTMath is highly challenging, with the best-performing model, o1-mini, solving only 32.57% of the problems, followed by o1-preview at 27.16%, and GPT-4o at 26.93%. Furthermore, we present the Reasoning-to-Coding of Thoughts (RCoT) approach, which encourages LLMs to engage in explicit reasoning prior to code generation, thereby facilitating the production of more sophisticated solutions and enhancing overall performance and efficiency. Furthermore, we also release the UTMath-Train training dataset (more than 70k samples), to support the community in further exploring mathematical reasoning. Our benchmark can be accessed via the following link: UTMath
@@ 1 Introduction
Figure 1: Comparison of UTMath with other benchmarks. On the left, GPT-4o easily solved one question but failed with a different numeric input. On the right, our benchmark is shown, where each problem includes multiple test cases, and a solution is correct only if all test cases are passed. We also propose a new prompting method RCoT in which the LLM first reasons through the problem and then generates code.
@@
However, classic benchmarks exhibit several limitations that impede the accurate and comprehensive assessment of these models’ capabilities Ahn et al. (2024). First, these benchmarks test models on narrowly defined problems with some specific numbers, which do not adequately assess generality to similar but varied scenarios as shown in Fig. 1. Second, their evaluation relies on predetermined rules or the method of LLM-as-a-Judge( Dubois et al. (2024); Zheng et al. (2023)) that usually failed with capricious responses of LLMs. For example, an accurate answer need to be extracted to exactly match the fianl answer in the dataset GSM8K, TheoremQA, and MATH dataset. While recent work has made great progress in developing new benchmarks, many of these approaches still fall short of addressing the fundamental limitations of earlier datasets. For instance, benchmarks like GSM-HARD Gao et al. (2023), GSM-IC Shi et al. (2023), GSM-Plus Li et al. (2024a), MetaMath Yu et al. (2023) have extended the dataset of GSM8K or MATH with some perturbation such as substitution, reversing, distractor insertion. These efforts, while valuable, are characterized by limited coverage and high costs. In this context, our work seeks to bridge these gaps by proposing a solid and robust benchmark that accurately evaluates the mathematical capabilities of LLMs.
@@
Drawing on evaluation methods from software development, we propose the design of a comprehensive set of unit tests for mathematical problems to rigorously assess the reasoning processes of LLMs. If a solution successfully passes all unit tests within a class of problems, it suggests that the reasoning underlying the solution is more reliable and trustworthy. Specifically, we introduce the UTMath, a novel benchmark derived from the On-Line Encyclopedia of Integer Sequences (OEIS) OEIS Foundation Inc. (2024). The benchmark consists of 1,053 cutting-edge problems spanning 9 mathematical domains, such as Number Theory and Geometry. Each problem is accompanied by more than 68 test cases that provide a set of inputs and their corresponding outputs.
@@
In terms of evaluation methodology, our benchmark requires models to derive a general solution for a class of problems, typically represented in the form of code. Compared to solving a problem defined by specific numbers, developing such a general solution is substantially more challenging, requiring higher levels of intelligence and reasoning ability. However, when we performed “Program of Thoughts (PoT)” Chen et al. (2022), wherethe model is required to perform reasoning and coding in a single response, it consistently tends to produce simpler and straightforward solutions. We surmise that this tendency may be influenced by the distribution of coding data. To address this, we introduced the “Reasoning-to-Coding of Thoughts (RCoT)”, which requires the LLM to perform mathematical reasoning in the first turn without any coding instruction, then writing code based on the reasoning. Compared to PoT, RCoT shifts the code distribution towards mathematics in the first turn, prompting more reasoning steps.
@@
UTMath is highly challenging,
We conducted a comprehensive study with 8 LLMs. Some of our key findings are summarized as follows: (1) with the best-performing model, o1-mini, solving only 32.57% of the problems, followed by o1-preview at 27.16%, and GPT-4o at 26.93%, these results demonstrate the difficulty of UTMath. (2) Modern LLMs perform poorly in Graph Theory, Group Theory, Geometry and Topology (Fig. 5). (3) With RCoT, all evaluated LLMs generated more efficient solutions, with most models achieving higher scores (Fig. 2). (4) RCoT can significantly improve the pass@k performance of LLMs (§ 5.4). (5) The quality of reasoning significantly impacts the accuracy and efficiency of the model’s final solution (§ 5.5). More interesting findings can be found in § 5. We hope our findings contribute to a deeper understanding of current reasoning ability of LLMs and the further development of models.
@@ 2.1 Benchmarks
Besides, researchers have also introduced many others focused on evaluating various aspects of LLMs like the robustness. These include GSM8K-based variants: GSM-8K-Adv  Anantheswaran et al. (2024), GSM-Hard  Gao et al. (2023), GSM-Plus  Li et al. (2024a), GSM-IC  Shi et al. (2023), and several independent benchmarks: RobustMath  Zhou et al. (2024b), MetaMathQA  Yu et al. (2023), PROBLEMATHIC  Anantheswaran et al. (2024), MATHCHECK  Zhou et al. (2024a), as well as other benchmarks  Li et al. (2024b, 2023).
The distinctions between our proposed benchmark and existing ones are as follows. (1) Multiple Case Validation. Instead of using single cases that can be memorized, our questions are sequence-based, allowing numerous cases for validating true understanding. (2) General Solutions. UTMath requires large models to solve problems by generating code, aiming for general solutions rather than problem-specific ones, reflecting a closer alignment with intelligence.
@@ 2.3 Prompting Methods
The Chain-of-Thought Wei et al. (2022) prompting technique encourages models to express reasoning steps in natural language before concluding. Similarly, the approach by  Kojima et al. (2022) uses the phrase "Let’s think step by step" to effectively guide large language models through their reasoning. Inspired by CoT, several effective prompting methods have been developed, such as Tree-of-Thoughts Yao et al. (2024), Graph-of-Thoughts Besta et al. (2024). . Program-of-Thought prompting Chen et al. (2022): PoT generates programs as the intermediate steps and integrates external tools like a Python interpreter for precise calculations, as well as other prompting methods Wang et al. (2023); Gao et al. (2023).
Our RCoT method stands out by dividing reasoning into two steps: reasoning and implementing based on reasoning. This segmentation provides deeper insights into LLM reasoning abilities. The advantages can be summarized as follows. (1) Enhanced Reasoning. Emphasizing reasoning allows large models to focus more on improving the quality of reasoning, thereby delivering higher-quality and more efficient solutions. (2) Modularity. By separating reasoning from implementation, the influence of coding on reasoning can be eliminated, providing a new paradigm for evaluating the reasoning ability through the code generated by the model.
Figure 3: UTMath generation pipeline. After downloading 23,238 Principle Sequences from OEIS and cleaning the data, 1,053 usable sequences were obtained. Descriptions were standardized by adding background information and improving readability (highlighted in green, also shown in Appendix B.2). Hard cases were introduced to enhance discriminative capability, including terms from later positions to prevent simplistic algorithms from passing.
@@ 3 UTMath Benchmark
The OEIS was established to document integer sequences of interest to both professional and amateur mathematicians, and it has become widely cited in the community. Most sequences are derived or updated from academic papers, contributing to their cutting-edge level of difficulty. As of February 2024, it contains over 370,000 sequences OEIS Foundation Inc. (2024). Each sequence is accompanied by an identification number, a brief description, some sample integers, links to relevant literature, and, where possible, program code for computing the sequences. An example sequence is shown in Appendix A.
3.2 Benchmark Construction.
UTMath is a cutting-edge and expansive benchmark designed to more accurately assess the mathematical reasoning abilities of LLMs. It consists of 1053 math problems, with each problem having an average of 68 test cases. The benchmark covers 9 mathematical domains, including not only common topics like number theory but also graph theory, group theory, topology, and geometry. The difficulty of UTMath is considered Cutting-Edge, with the majority of the sequences that form the problems having been studied in academic papers. UTMath was obtained as follow (see also Fig.3).
Data Crawling.
@@
OEIS provides users with a list of principal sequences  [^1] https://oeis.org/wiki/Index_to_OEIS, which are most important sequences defined by OEIS. OEIS categorizes these sequences into sections based on the first 2-3 letters of their content themes. By scraping the category tags within each section and the AIDs of their subordinate sequences, we obtained 23,238 principal sequences’ AIDs. OEIS provides an interface to request the JSON data of the HTML page for each sequence using its AID  [^2] https://oeis.org/wiki/JSON_Format. By passing the sequence AIDs to this interface, we acquired the JSON data for these 23,238 sequences.
Data Cleaning.
@@
We found that some of the sequences we collected did not meet our criteria and should be removed, with further details provided in the Appendix B. Here are several main situations.
@@
$\bullet$ Hard to solve, few terms are discoverable. A portion of the sequences retrieved are marked as “hard” in the keyword field of their entries in OEIS. According to OEIS, “Any sequence which can be extended only by new ideas, rather than more computation deserves keyword: hard. Similarly, if computing a term of the sequence would probably merit a paper in a peer-reviewed journal (discussing the result, the algorithm, etc.)”  [^3] https://oeis.org/wiki/User:Charles_R_Greathouse_IV/Keywords/difficulty Another related keyword attribute is "fin" (finite), indicating sequences with limited length. For our purposes, sequences should be infinitely derivable.
@@
$\bullet$ Difficult to Generate Programmatically. In OEIS, most sequences are provided with fields such as Mathematica, program, or formula, but not all sequences include these details. We assume that the sequences without these fields may be difficult to generate programmatically.
@@
$\bullet$ Simple Sequences. Some sequences are too simple to require any reasoning. We use GPT-4o to determine if a sequence requires reasoning or just implementation; if mostly implementation, it’s excluded. For instance, A000178444https://oeis.org/A000178: ’Superfactorials: product of the first n factorials,’ a sequence requiring only implementation, will be excluded.
After addressing the aforementioned issues, we ultimately obtained 1053 sequences.
Dataset Size Level Multi-test Efficiency Metric Output
College Math 2,818 University Accuracy Text
GSM8K 1,319 Elementary school Accuracy Text
MATH 5,000 High school Accuracy Text
RobustMath 300 High school Accuracy Text
OlympiadBench 8,476 Competition Accuracy Text
TheoremQA 800 University Accuracy Text
UTMath(ours) 1,053 Cutting-edge Pass Rate Code
Table 1: Comparison between UTMath and other benchmarks. UTMath offers a cutting-edge benchmark with a comprehensive set of 1,053 problems across multiple mathematical domains, providing a more accurate evaluation of LLMs’ mathematical reasoning capabilities.
Standardization of Question Statements.
As a specialized academic database in the field of mathematics, OEIS provides a wealth of useful information for each sequence. However, we have found that some sequences cannot be directly used with the descriptions provided by OEIS as problem statements, primarily for the following reasons: (1) Specialized Terminology. Some sequence descriptions use complex math terms that need examples or explanations to be clear. Using them directly as problems might test mathematical knowledge rather than reasoning skills. So, it is important to explain key concepts to focus on reasoning and reduce the extra knowledge needed.(2) Brevity and Ambiguity. Some sequence descriptions are excessively brief and lack a clear definition of what a(n) is. We used GPT-4o to standardize these by adding background info and making the language smoother. The prompts we used are provided in the Appendix C and an example is shown in Appendix B.2. Despite our requirement that ensuring the consistency of meaning between the original and processed descriptions, hallucinations can still occur. To mitigate this, we performed a manual verification of the standardized problem statements to ensure they matched the original descriptions in meaning and were easy to comprehend.
Hard Test Cases Mining.
@@
The primary goal of this paper is to evaluate the reasoning ability of LLMs. Generally, more efficient solutions to a problem imply stronger reasoning capabilities. Therefore, we aim for our evaluation to distinguish whether a solution is efficient. However, in the OEIS, each sequence only lists the first few n terms, normally n<100, which can be obtained without requiring particularly efficient methods. This limitation prevents the evaluation from effectively distinguishing between efficient and inefficient solutions. An obvious fact is that the difficulty of computing the first 10 terms of a sequence within a time limit is significantly different from computing terms starting from the $10^{6}$th term. Therefore, we aim to create more challenging test data to better assess the reasoning capabilities of LLMs.
@@
Fortunately, many OEIS sequences include corresponding Mathematica code that can be regarded as the ground-truth solution for each problem. We extract this Mathematica code for each sequence, formalizing it to compute the first $N$ terms, $A_{1},...,A_{N}$, of the sequence. We determine the maximum value of $N_{max}$ for which the code can compute the sequence within 10 seconds, where we set $10^{6}$ as the upper bound. Finally, we add the last 10 terms $A_{N_{max-9}},...,A_{N_{max}}$ into our benchmark as the hard test cases to evaluate the complexity of a solution. Our experiments demonstrate that these cases precisely differentiate more efficient and intelligent solutions.